OpenStudy (yinshy):

A cylindrical water tank has a diameter of 6 ft and a height of 8 ft. What is the exact volume of the tank? A. 24π ft³ B. 48π ft³ C. 72π ft³ D. 288π ft³

OpenStudy (anonymous):

it's C

OpenStudy (yinshy):

Which measure is closest to the volume of the cylinder?

OpenStudy (anonymous):

\[72\pi ft\]

OpenStudy (anonymous):

226ft to the nearest whole number

OpenStudy (rushda99):

\[Volume = \pi r^2 h\] = π x 3^2 x 8 = 72π ft

OpenStudy (anonymous):

C.